OpenStudy (anonymous):

Solve linear diff. Equation (Xlogx)dy/dx +y = 3x^3

OpenStudy (accessdenied):

I am assuming that log x is the natural log here?

OpenStudy (anonymous):

\[xln(x) \frac{dy}{dx}+y=3x^3\] \[xln(x) \frac{dy}{dx} f(x)+yf(x)=f(x)3x^3\] Define \[xln(x) f(x)=\frac{df(x)}{dx}\] Use this to find explicit formula for f(x) Then, \[\frac{dy}{dx} f'(x)+yf(x)=f(x)3x^3\] turns into \[\frac{d(y*f(x))}{dx}=f(x)3x^3\]

OpenStudy (anonymous):

\[yf(x)=\int\limits f(x)3x^3dx\]Divide by f(x)

OpenStudy (anonymous):

All you have to do now is (1) Understand what I did- ask if confused (2) Find explicitly the formula for f(x) and solve for y

OpenStudy (anonymous):

Sorry bout my weakness but i dont understand what is f(x)

OpenStudy (anonymous):

I use integrating factor to solve this so i am not able to get e^integral(1/xlogx)

OpenStudy (accessdenied):

I believe with the aforementioned method by henpen, \(f(x)\) is acting as our integrating factor. If we want to figure out the integrating factor using your formula: \(\displaystyle \large{ e^{\int \frac{1}{x \log x} \; \text{d}x} } \) We just need to figure out this integral: \(\int \frac{1}{x \log x} \; \text{d}x \) Which we should notice the log x and a 1/x in the integral. That looks good for u-substitution with u=log x to figure out. We can also solve henpen's equation, which should also work out. :)

OpenStudy (anonymous):

Thnx

OpenStudy (accessdenied):

You're welcome!

OpenStudy (anonymous):

Is ther an ipad app for this amazing site?

OpenStudy (accessdenied):

Hm, I'm not sure there is, although I have heard talk about some people wanting one.